OpenStudy (anonymous):

If I want to know what is 20200/e(0.07695)(16), how do I input the "e" in the calculator?

OpenStudy (anonymous):

I understand the e = ln but Im not getting the same anwer my instructors gets

OpenStudy (anonymous):

the problem is A=pe^rt, where A=20200/e(0.076950(16) his answer is 5897.23

OpenStudy (anonymous):

The r =7.695% t= 16 years) or 16. I need help with finding e, please

OpenStudy (anonymous):

Hi, do you have any idea how to solve this?

OpenStudy (nowhereman):

e is Euler's constant at about 2.71828. A good calculator should have it stored. It is the following limit: \[e = \lim_{n→∞}{(1+1/n)^n} = \sum_{n=0}^∞{\frac{1}{n!}}\]\[ = 1 + 1\left(1+\frac 1 2 \left(1 + \frac 1 3\left(1+\frac 1 4\left(1+\frac 1 5\left(1 + \cdots\right)\right)\right)\right)\right)\]

OpenStudy (anonymous):

That looks like stats. Would the same be applied to basic algebra?

OpenStudy (nowhereman):

It has nothing to do with statistics. Euler's constant is one of the most basic (next to π) constants of mathematics. It occurs very often, especially in analysis.

OpenStudy (anonymous):

That method is not getting me the same answer as my instructor

OpenStudy (nowhereman):

Well, the only way I can get a nearly equal is by calculating \[\frac {20200}{1.08^{16}}\] But that is not really accurate if you are talking about interests. Then I would do \[\frac{20200}{1.07695^{16}} = 6169.11\]

OpenStudy (anonymous):

Here's the entire question: what principal should be deposited at 7.695% to ensure the acct will be worth $20200 in 16 years. I have 2020/e(0.07695)(16)

OpenStudy (anonymous):

The formula is A=pe^rt. Im not understanding "e" and its throwing my answer off. His answer is $5897.23

OpenStudy (anonymous):

In this case do you know how to calculate the e?

OpenStudy (nowhereman):

That formula looks strange. Lets say initially you deposit a, and you have an interest rate of r = 7.695% = 0.07695 then after the first year you have your initial amount PLUS the interest: \[a_1 = a + r\cdot a = a\cdot (1 + r)\] and one year later the same is true based on a1. So after n years you get \[a_n = a\cdot(1+r)^{n}\] No e there.

OpenStudy (anonymous):

Do you know how he got $5897.23?

OpenStudy (nowhereman):

I told you above. He rounded the interest rate to 8% and used the correct formula \[a = \frac{a_n}{(1+r)^n}\] at least that is the closest I can come to that solution.

OpenStudy (anonymous):

In this case A = 20200/1+16?

OpenStudy (anonymous):

20200/(1+(0.07695)(16)?

OpenStudy (anonymous):

You have to use \(A = Pe^{rt}\) When doing continuously compounded interest. e is the base of the exponential function. You should have a button on your calculator for it (typically it's the same button as the LN button). And if not you can use an approximate value of 2.71828183